# Conversational AI increases political knowledge as effectively as self-directed internet search

### v2

@@ Abstract
Conversational AI systems are increasingly being used in place of traditional search engines to help users complete information-seeking tasks. This has raised concerns in the political domain, where biased or hallucinated outputs could misinform voters or distort public opinion. However, in spite of these concerns, the extent to which conversational AI is used for political information-seeking, as well the potential impact of this use on users’ political knowledge, remains uncertain. Here, we address these questions: First, in a representative national survey of the UK public (N = 2,499), we find that in the week before the 2024 election as many as 32% of chatbot users—and 13% of eligible UK voters—have used conversational AI to seek political information relevant to their electoral choice. Second, in a series of randomised controlled trials (N = 2,858 total) we find that across issues, models, and prompting strategies, conversations with AI increase political knowledge (increase belief in true information and decrease belief in misinformation) to the same extent as self-directed internet search. Taken together, our results suggest that although people in the UK are increasingly turning to conversational AI for information about politics, this shift may not lead to increased public belief in political misinformation.
@@ Introduction
Conversational AI systems (or chatbots) such as ChatGPT, Claude and Gemini are now used regularly by hundreds of millions of people across the world. Analysis of consumer usage reveals that among the most common use cases is the seeking of information for private and professional purposes, including both general knowledge and practical advice [Anthropic, 2024]. As public usage of chatbots continues to increase, a particular concern is that LLMs will produce unreliable or biased answers when queried about current affairs or political issues. This could be damaging to democracy, which relies on the electorate having access to reliable information, especially before elections [Summerfield et al., 2024].
@@
However, two main uncertainties make the extent of this risk unclear. First, it is unclear whether people trust AI enough to use it for political information-seeking in high-stakes political moments, such as in the lead-up to national elections. Among the general public, trust in AI is low: surveys consistently report that a majority of people do not trust AI to be used as tools to provide medical care, legal advice, or to assist human journalists [Newman et al., 2024; Gillespie et al., 2023; McClain, 2024]. One recent survey showed that levels of trust in AI were comparable to those for politicians, advertising executives, and social media influencers who are themselves the least trusted of all professions [Laher, 2024]. Given this, it is an open question whether people will turn to conversational AI for political guidance during elections, even when such tools are available.
@@
Second, it is unclear whether conversational AI systems present functional substitutes for traditional search engines on common political issues. On the one hand, it is widely acknowledged that AI models have a problem with factuality (AI developers have devoted considerable energy to studying what they call ‘hallucinations’ in AI models [Huang et al., 2024; Ji et al., 2023; Augenstein et al., 2024; Wang et al., 2023]) and there has been concern that models may be politically biased, especially in a progressive or libertarian direction [Santurkar et al., 2023; Hartmann et al., 2023; Röttger et al., 2024]. On the other hand, significant progress has been made in tackling these risks: developers have implemented techniques such as retrieval-augmented generation (RAG) [Lewis et al., 2020], knowledge-graphs [Agrawal et al., 2024], fine-tuning on factuality rankings [Tian et al., 2024], and semantic entropy methods for detecting confabulations [Farquhar et al., 2024]. As a result, it remains unclear if reliability issues would substantively impact users’ political knowledge in a real-world information-seeking setting.
@@
Here, we address both of these uncertainties. First, in a representative survey one week after the UK 2024 general election, we find that as many as 13% of eligible voters may have used conversational AI to find information relevant to their electoral choice. Second, in an RCT we find that participants who used a chatbot (GPT-4, Claude, or Mistral) to research factual information related to issues of concern for UK voters increased their political knowledge to the same extent as participants who researched the same issues using an internet search engine. In follow-up RCTs, we find that even when LLMs were explicitly prompted to use sycophantic or persuasive techniques, participants’ belief formation did not differ from those interacting with standard unprompted AI models.
@@ Discussion
We demonstrate that conversational AI now sits alongside internet search as a commonly used source of information during high-stakes political moments like national elections. Moreover, across four salient issues and three model families, researching with chatbots raised belief in true facts and lowered belief in false claims to the same extent as self-directed internet search, even when the chatbots were prompted to be persuasive or sycophantic. These findings stand in contrast to the popular assumption that the use of chatbots for seeking information about news or current affairs may inherently erode political knowledge.
This finding has two main implications. First, our results suggest that the large body of prior work suggesting low levels of public trust in AI belies actual public usage for information-seeking: in fact, our data suggest that information-seeking is the most popular use case for chatbots among UK citizens (surpassing professional use, writing/translation, and practical advice), and that that users found their chatbots to be useful, accurate, and un-biased. Second, our experiment suggests that contrary to widespread concern about AI reliability and hallucinations, models do not increase belief in false information when used to research current affairs and political issues. These results could suggest that developers’ current guardrails may be sufficient to keep average informational effects neutral, and could open space to refine AI for rapid, reliable public learning during high-stakes events. Chatbot users were also able to research more quickly compared to search, an efficiency incentive which could portend further adoption of chatbots.
@@
Within the above limits, our results suggest that for everyday information-seeking, today’s chatbots may perform on par with self-directed internet search—potentially with no cost to political knowledge.
@@ Methods
Issue agreement and extremism data before and after search/AI conversation were analyzed using three Bayesian multilevel GLMs [Luettgau et al., 2025; Dubois et al., 2025]. We specified GLMs with ordered-logistic likelihood functions to model the ordinal categories of issue agreement responses. For extremism data, we defined GLMs with Binomial likelihood function. GLMs were fitted using sampling-based Bayesian inference using Numpyro [Phan et al., 2019] for Markov Chain Monte Carlo sampling (using No-U-Turn-Sampler – NUTS – a variant of Hamiltonian Monte Carlo [Hoffman and Gelman, 2011]) to estimate the posterior distribution of linear model parameters. These parameters include different intercept and slope parameters that combine linearly to influence the likelihood of ordinal or Binomial responses.
@@
For GLM comparisons and to identify the best-fitting GLM for the observed data, we used the Widely Applicable Information Criterion (WAIC [Watanabe, 2010]). Parameter estimates were considered non-zero if the Highest Posterior Density Interval (HPDI) around the parameter did not contain zero. The HPDI was compared to a region of practical equivalence (ROPE), i.e., an interval of parameter values [0.05;0.05][-0.05;0.05] representing the null hypothesis of the parameter being equivalent to 0.

### v3

@@ Abstract
Conversational AI systems are increasingly being used in place of traditional search engines to help users complete information-seeking tasks. This has raised concerns in the political domain, where biased or hallucinated outputs could misinform voters or distort public opinion. However, in spite of these concerns, the extent to which conversational AI is used for political information-seeking, as well the potential impact of this use on users’ political knowledge, remains uncertain. Here, we address these questions: First, in a representative national survey of the UK public (N = 2,499), we find that in the week before the 2024 election as many as 32% of chatbot users – and 13% of eligible UK voters – have used conversational AI to seek political information relevant to their electoral choice. Second, in a series of randomised controlled trials (N = 2,858 total) we find that across issues, models, and prompting strategies, conversations with AI increase political knowledge (increase belief in true information and decrease belief in misinformation) to the same extent as self-directed internet search. Taken together, our results suggest that people in the UK are increasingly turning to conversational AI for information about politics. These findings substantially extend prior work by demonstrating that conversational AI’s effects on political knowledge generalise across multiple topics, political perspectives, and model families, suggesting that the shift toward AI-assisted political information-seeking may not lead to increased public belief in political misinformation.
@@ Introduction
Conversational AI systems (or chatbots) such as ChatGPT, Claude and Gemini are now used regularly by hundreds of millions of people across the world. Analysis of consumer usage reveals that among the most common use cases is the seeking of information for private and professional purposes, including both general knowledge and practical advice [anthropic2024clio]. As public usage of chatbots continues to increase, a particular concern is that LLMs will produce unreliable or biased answers when queried about current affairs or political issues. This could be damaging to democracy, which relies on the electorate having access to reliable information, especially before elections [summerfield2024democracy].
@@
However, two main uncertainties make the extent of this risk unclear. First, it is unclear whether people trust AI enough to use it for political information-seeking in high-stakes political moments, such as in the lead-up to national elections. Among the general public, trust in AI is low: surveys consistently report that a majority of people do not trust AI to be used as tools to provide medical care, legal advice, or to assist human journalists [newman2024digital; gillespie2023trust; mcclain2024chatgpt]. One recent survey showed that levels of trust in AI were comparable to those for politicians, advertising executives, and social media influencers who are themselves the least trusted of all professions [laher2024trust]. Given this, it is an open question whether people will turn to conversational AI for political guidance during elections, even when such tools are available.
@@
Second, it is unclear whether conversational AI systems present functional substitutes for traditional search engines on common political issues. On the one hand, it is widely acknowledged that AI models have a problem with factuality (AI developers have devoted considerable energy to studying what they call ‘hallucinations’ in AI models [huang2024hallucination; ji2023hallucination; augenstein2024factuality; wang2023factuality]) and there has been concern that models may be politically biased, especially in a progressive or libertarian direction [santurkar2023opinions; hartmann2023ideology; rottger2024compass]. On the other hand, significant progress has been made in tackling these risks: developers have implemented techniques such as retrieval-augmented generation (RAG) [lewis2020retrieval], knowledge-graphs [agrawal-etal-2024-knowledge], fine-tuning on factuality rankings [tian2024finetuning], and semantic entropy methods for detecting confabulations [farquhar2024detecting]. As a result, it remains unclear if reliability issues would substantively impact users’ political knowledge in a real-world information-seeking setting.
Here, we address both of these uncertainties with a comprehensive, multi-dimensional examination of conversational AI’s effects on political knowledge. Recent work [taylor2024ai] provided initial evidence that conversational AI can improve factual knowledge on specific political topics, examining the effects on beliefs about crime rates and vaccine efficacy. However, this preliminary work was limited to single-issue investigations favouring liberal-aligned factual information, without built-in controls for general knowledge acquisition or examination of effects across the political spectrum. Here, we substantially extend this work by: (1) examining multiple political topics simultaneously within subjects to assess generalisability; (2) including politically balanced true/false information representing both progressive and conservative viewpoints; (3) implementing within-subject controls for non-researched topics to isolate AI-specific effects; and (4) testing multiple models, prompting strategies, and secondary outcomes including trust, private beliefs, and extremism. Additionally, we provide the first nationally representative data on actual chatbot usage for political information-seeking during a national election.
@@
In the present study, we first show in a representative survey one week after the UK 2024 general election that as many as 13% of eligible voters may have used conversational AI to find information relevant to their electoral choice. Second, in an RCT we find that participants who used a chatbot (GPT-4, Claude, or Mistral) to research factual information related to issues of concern for UK voters increased their political knowledge to the same extent as participants who researched the same issues using an internet search engine. In follow-up RCTs, we find that even when LLMs were explicitly prompted to use sycophantic or persuasive techniques, participants’ belief formation did not differ from those interacting with standard unprompted AI models.
@@
Our work makes three key contributions: (1) Replication: We replicate recent findings that conversational AI can improve political knowledge on specific issues [taylor2024ai]; (2) Generalisation: We demonstrate that these effects generalise across multiple political topics, model families, prompting strategies, and remain consistent across politically balanced information representing both progressive and conservative viewpoints; (3) Comprehensive assessment: We provide the first assessment of AI’s effects on broader epistemic health indicators (trust, private beliefs, extremism) and the first nationally representative data on real-world chatbot usage for political information during a national election.
@@ Results
Our results replicate recent findings that conversational AI can improve factual political knowledge on specific issues [taylor2024ai]. We substantially extend this work in several key ways. First, while prior work examined single issues in isolation, our within-subjects design demonstrates that these effects generalize across four politically salient topics simultaneously (climate change, immigration, criminal justice, COVID-19). Second, our inclusion of both progressive- and conservative-aligned true information (see Supplementary Information) demonstrates that AI’s knowledge-enhancing effects are not limited to liberal-favored facts. Third, our within-subject control topics (non-researched issues) allow us to rule out generic effects of questionnaire repetition or participant engagement, isolating AI-specific learning effects. Fourth, we extend beyond single knowledge questions to examine broader epistemic health indicators including trust, private political beliefs, and extremism – none of which showed differential effects between AI and traditional search.
@@ Discussion
Replicating and extending recent work [taylor2024ai], we demonstrate that conversational AI now sits alongside internet search as a commonly used source of information during high-stakes political moments like national elections. Going beyond prior single-issue investigations, we show that across four salient issues, politically balanced true and false information, and three model families, researching with chatbots raised belief in true facts and lowered belief in false claims to the same extent as self-guided internet search, even when the chatbots were prompted to be persuasive or sycophantic. Our within-subject controls for non-researched topics and secondary outcomes (trust, private beliefs, extremism) further demonstrate that these effects represent genuine knowledge acquisition rather than generic engagement or politically-biased persuasion. These findings stand in contrast to the popular assumption that the use of chatbots for seeking information about news or current affairs may inherently erode political knowledge.
Our comprehensive, multi-topic examination has three main implications. First, we replicate prior work showing that AI can improve political knowledge, while demonstrating for the first time that this effect generalizes across multiple topics, political viewpoints, and is not limited to liberal-favored information. Second, our results suggest that the large body of prior work suggesting low levels of public trust in AI belies actual public usage for information-seeking: in fact, our data suggest that information-seeking is the most popular use case for chatbots among UK citizens (surpassing professional use, writing/translation, and practical advice), and that users found their chatbots to be useful, accurate, and unbiased. Third, our experiment suggests that contrary to widespread concern about AI reliability and hallucinations, models do not increase belief in false information when used to research current affairs and political issues across diverse topics and political perspectives.
@@
While our study replicates the core finding of regarding AI’s capacity to improve factual knowledge [taylor2024ai], our multi-topic, politically-balanced design with built-in controls provides substantially stronger evidence for the generalisability and political neutrality of these effects. However, like prior work, we remain limited to examining factual questions with objectively verifiable answers, and future research should examine more subjective or contested political claims where the ’correct’ answer is less clear-cut.
Within the above limits, our results suggest that for everyday information-seeking, today’s chatbots may perform on par with self-directed internet search –potentially with no cost to political knowledge.
@@ Methods
Issue agreement and extremism data before and after search/AI conversation were analyzed using three Bayesian multilevel GLMs [luettgau2025hibayeshierarchicalbayesianmodeling; dubois2025skewedscorestatisticalframework]. We specified GLMs with ordered-logistic likelihood functions to model the ordinal categories of issue agreement responses. For extremism data, we defined GLMs with Binomial likelihood function. GLMs were fitted using sampling-based Bayesian inference using Numpyro [Phan2019] for Markov Chain Monte Carlo sampling (using No-U-Turn-Sampler – NUTS – a variant of Hamiltonian Monte Carlo [Hoffman2011]) to estimate the posterior distribution of linear model parameters. These parameters include different intercept and slope parameters that combine linearly to influence the likelihood of ordinal or Binomial responses.
@@
For GLM comparisons and to identify the best-fitting GLM for the observed data, we used the Widely Applicable Information Criterion (WAIC [WatanabeSWATANAB2010]). Parameter estimates were considered non-zero if the Highest Posterior Density Interval (HPDI) around the parameter did not contain zero. The HPDI was compared to a region of practical equivalence (ROPE), i.e., an interval of parameter values [0.05;0.05][-0.05;0.05] representing the null hypothesis of the parameter being equivalent to 0.